Find an algebraic expression. The perimeter of a rectangle, given that the length is x yards and the width is 10 yards shorter.
Found 2 solutions by drj, butofcourseric:
Answer by drj(1380)

Find an algebraic expression. The perimeter of a rectangle, given that the length is x yards and the width is 10 yards shorter.

Step 1. Perimeter P means adding up all four sides of the rectangle.

Step 2. x-10 be the width of the rectangle since the width is 10 yards shorter.

Step 3. P=x+x+x-10+x-10=4x-20

Step 4. ANSWER: Perimeter P=4x-20 (in yards).

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

And good luck in your studies!

Respectfully,
Dr J
http://www.FreedomUniversity.TV

Answer by butofcourseric(2)

P = x+x+x-10+x-10
Ans: P= 4x-20